OpenStudy (anonymous):

4a-14=4a-14 show me how to solve for a?

OpenStudy (anonymous):

well you would add 14+4 and get the answer

OpenStudy (anonymous):

so the answer would be 18

OpenStudy (anonymous):

does that help

OpenStudy (anonymous):

the answer is always true. 4a-14 does equal 4a-14

OpenStudy (anonymous):

it is equal because it is the same problem

OpenStudy (anonymous):

I just need to figure out how to solve A, test in 20 minutes

OpenStudy (anonymous):

correct me if im wrong but im sure this problem has no solution

OpenStudy (anonymous):

the answer is givin within the equation

OpenStudy (anonymous):

but he is having to solve for a

OpenStudy (anonymous):

he has to work the equation out to find the answer

OpenStudy (anonymous):

4a-14 = 4a-14 +14 +14 4a = 4a -4 -4 0 = 0 No solution